# SAEs Are Good for Steering – If You Select the Right Features

### v1

@@ Abstract
Dana Arad1thanks: Work partially done during an internship at Amazon.  Aaron Mueller1,2  Yonatan Belinkov1 1Technion – Israel Institute of Technology  2Northeastern University {danaarad@campus., belinkov@ }technion.ac.il aa.mueller@northeastern.edu
@@ 1 Introduction
Sparse autoencoders (SAEs) have shown promise in extracting human-interpretable features from the hidden states of language models (LMs) Bricken et al. (2023); Templeton et al. (2024). One appealing usage of SAEs is to enable fine-grained interventions such as generation steering O’Brien et al. (2024); Durmus et al. (2024); Marks et al. (2025). However, selecting the right features for intervention is an open problem. Current approaches typically select features to steer based on their activation patterns, i.e., the input texts that most strongly activate a given feature Huben et al. (2024). While input-based activations can reveal meaningful patterns, recent work highlights a critical limitation: a feature’s activations are not necessarily the same as its causal effect on the model’s output Durmus et al. (2024); Paulo et al. (2024); Gur-Arieh et al. (2025). As a result, the way features are selected can lead to suboptimal steering, reducing its consistency and reliability Durmus et al. (2024).
@@
In this work, we formalize two distinct roles that features can play: input features, which capture patterns within the model’s input, and output features, whose main role is to directly influence the tokens the model generates. To find them, we propose input scores and output scores. First, we obtain a representative set of tokens for each feature by applying the logit lens to SAE weights; this projects the weights directly into the vocabulary space nostalgebraist (2020); Bloom and Lin (2024). We define input features as having high input scores—i.e., high overlap between their top-activating tokens and top logit lens tokens. We define output features as having high output scores—i.e., intervening on the feature increases the probability of its top logit lens tokens in the final output distribution. Notably, the input score can be computed in parallel for all features over a general dataset; the output score requires only one forward pass and no concept-specific data. We quantitatively show that these roles rarely co-occur and tend to emerge at different layers in the model. Specifically, features in earlier layers primarily act as detectors of input patterns, while features in later layers are more likely to drive the model’s outputs, consistent with prior analyses of LLM neuron functionality Lad et al. (2024); Marks et al. (2025).
@@
By calculating the input and output scores of features extracted from Gemma2 (2B and 9B) Team et al. (2024), we show that features with high output scores are more effective for coherent and high-quality steering (Figure 1 demonstrates the feature selection process). Figure 2 demonstrates the difference when steering with input features vs. output features: steering with a feature that has a high input score but low output score fails to meaningfully influence the generation. In contrast, steering using a feature with high output score and low input score yields better steering, as well as more fluent and semantically coherent completions.
@@
We demonstrate the effectiveness of these insights on the recent AxBench Wu et al. (2025), a benchmark for evaluating steering methods. While AxBench found SAEs to be poor for steering, our feature selection results in a 2–3x improvement, causing SAE steering (an unsupervised method) to score significantly closer to supervised methods like LoRA Xu et al. (2024).
@@
  • We propose an efficient method for finding features effective for steering.
@@ 2.1 Sparse Autoencoders
Sparse Autoencoders (SAEs) were recently proposed as a method to address the problem of polysemanticity, where individual neurons entangle multiple unrelated concepts, and where a single concept may be distributed across many neurons Bricken et al. (2023). Given a hidden representation $x\in\mathbb{R}^{n}$, an SAE consists of an encoder and a decoder, defined as:
@@
The encoder maps the latent $x$ into a higher-dimensional sparse vector $a(x)$, or $a$ for short, which we refer to as the activations. The decoder reconstructs $x$ from $a$ as a sparse linear combination of the learned features, given by the columns of $W_{\text{dec}}$. Sparsity and non-negativity of the activations are enforced through the non-linearity $\sigma$, often JumpReLU Rajamanoharan et al. (2024), and regularization.
@@ 2.2 The Logit Lens
The logit lens is a widely used interpretability tool for analyzing the hidden representations of language models nostalgebraist (2020). Given a hidden state $x\in\mathbb{R}^{n}$ at any layer of the model, the logit lens passes $x$ through the final layer norm, LN, then projects $x$ onto the vocabulary space by applying the unembedding matrix $W_{\text{unembed}}\in\mathbb{R}^{n\times|\mathcal{V}|}$, where $\mathcal{V}$ is the model’s vocabulary. This produces a vector of predicted logits:
@@
Recent work has demonstrated that the logit lens can be applied not only to hidden representations, but also model weights Dar et al. (2023), gradients Katz et al. (2024), and even in multi-modal settings Toker et al. (2024).
Bloom and Lin (2024) suggested applying the logit lens to SAE feature weights as a way to interpret their roles. To interpret an SAE feature $f_{i}$, corresponding to $W^{i}_{\text{dec}}$, the $i$-th column in the decoder matrix, we compute:
@@ 2.3 Steering LMs
We define steering as influencing the output of an LM towards a desired concept. Successful steering should maintain the quality and coherency of the generated text. In other words, we seek a minimal change to a model’s computation that adds or subtracts a concept’s influence. Steering can be done by various methods, recently using SAEs Durmus et al. (2024); Wu et al. (2025).
Formally, given a model $M$ and a prompt $x$, we obtain a steered text $\hat{y}$ by applying an intervention $\Phi(\cdot)$ on some intermediate representation $h$:
$\hat{y}=M_{h\leftarrow\Phi(h)}(x)$ (5)
Similarly to Templeton et al. (2024), in order to steer an LM towards a concept $c$ encoded in an SAE feature $f_{i}$ at layer $l$, we define $\Phi$ as follows: first, we pass a prompt prefix $p$ through the model. At layer $l$, we pass the latent representation $x^{l}$ through the SAE encoder to obtain the activations vector, $a$. We record the max-activating feature, denoted $a_{\text{max}}$. Then, we obtain a new activation vector using steering factor $s$:
@@
Similarly to Aleksandar Makelov (2024), to evaluate steering success we measure the generation success w.r.t. $\ell_{k}$ by calculating the number of appearances of any token in $\ell_{k}$ in the generated text. Given a set of sentences $S$:
[…]
Additionally, we use perplexity measured using Gemma2-9B to quantify the generation coherence of the entire generated text.
@@ 3.1 Input Features
In practice, we use the pre-computed activations from Neuronpedia Lin (2023) to obtain the sentences $S$. We verify that $S$ has at least 20 sentences and take a maximum of 100 sentences per feature.
@@ 3.2 Output Features
Output features were first mentioned by Paulo et al. (2024) as features whose effect on the model’s output can be easily explained in natural language. Natural language explanations predispose us to errors in both precision and recall (Huang et al., 2023); therefore, given a target concept, we quantify steering quality as consistency between the set of logit lens tokens and the set of tokens that the steering operation promotes.
@@
To measure the effect of a feature on the model’s output distribution, we perform an intervention during a forward pass and evaluate the change in the rank and probability assigned to tokens in $\ell$. We first use a neutral prompt ($x=$ ”In my experience,”) to obtain the model’s prior distribution over token ranks and probabilities. Then, we intervene on the feature’s activation value using a large steering factor (we use $10$), as in Equation 6.
@@
where $V$ is the model’s vocabulary. If $x$ is a neutral prompt, then $S_{out}\propto P(\mathcal{M}_{h\leftarrow\Phi(h)})$, so we can compute $S_{out}$ quickly using only a single forward pass by only computing the rank-weighted probability after the intervention.
@@ 3.3 Experimental Setup
We focus our analysis on Gemma-2 (2B and 9B), for which high-quality SAEs are available in Gemma-Scope Team et al. (2024); Lieberum et al. (2024). We use the 16k SAEs. Our analysis includes 100 features randomly sampled from each layer (2,600 and 4,200 features in total for Gemma-2 2B and 9B, respectively).
(a) Gemma2-2B.
(b) Gemma2-9B.
Figure 3: Input and output scores across layers in Gemma2-2B and Gemma2-9B. The solid lines represent the median input score (blue) and output score (magenta), while the shaded regions denote the interquartile range (25th to 75th percentile), capturing the variability across features within each layer. Early layers are characterized by features with high input scores, while high output scores emerge in later layers.
@@ 3.4 Results
Figure 3 shows the distribution of input and output scores across layers for Gemma2-2B and Gemma2-9B. In early layers (0–50% of model depth), features tend to have high input scores and near-zero output scores, suggesting they are predominantly input-aligned. Later layers (66–100% of model depth) show an opposite trend: input scores drop to near-zero, while output scores increase significantly. These later-layer features no longer reflect the tokens they are activated on, but instead align with the tokens they promote in the model’s output, indicating a shift toward output-aligned behavior. Interestingly, middle layers exhibit low scores for both metrics, suggesting that these features may play intermediate roles that are neither purely input-aligned nor strongly output-promoting.
A natural concern is that the logit lens tends to be uninterpretable in early layers nostalgebraist (2020). Empirically, we find that many features promote a coherent and human-understandable set of tokens, as reflected by high input scores as early as layers 0 and 1. (See Appendix B for examples.)
@@ 4.1 Experimental Setup
We use 50 prompt prefixes and generate up to 20 tokens, obtaining 50 generated texts for each feature (more details in Appendix C). For each feature we calculate the mean generation success across the generated texts, and filter out steering factors leading to generation success greater than $3$. Intuitively, the generation success measures the rate in which the model generates concept-related tokens. Based on early experiments, we find that 3 is a good upper value that balances steering and coherence. We choose the optimal steering factor as the one that maximizes Gen $\frac{\text{Gen Success@20}}{\text{Perplexity}}$, where we normalize both metrics to a 001111 range by dividing with the maximum value across all data samples.
@@ 4.2 Qualitative Results
Figure 2 demonstrates steering with two features: one having a high input score and low output score (an input feature), and the other, an output feature, having a high output score and low input score. Steering using the input feature fails to meaningfully influence the generation. When using a high steering factor, this results in repeatedly generating tokens related to the feature’s activation, as if continuing from the word ”primary”. In contrast, steering using a feature with high output score and low input score using an optimal steering factor yields fluent and semantically coherent completions. Additional examples are shown in Appendix C.
(a) Gemma2-2B.
(b) Gemma2-9B.
Figure 5: Even in later layers of the model (16–25 for Gemma2-2B and 24–41 for Gemma2-9B), filtering features with low output scores increases mean generation success. Magenta: filtering by output scores. Green: filtering random sets of features of the same size.
@@ 4.3 Quantitative Results
Figure 4 shows the mean generation success @ 20 of steered generations when filtering out features with output scores below varying thresholds (magenta). As the threshold increases, performance improves steadily, indicating that features with higher output scores consistently lead to more successful steering. We observe an increase in the mean generation success score from around $0.5-0.6$ for both models without any filtering, to 1.11.11.11.11.41.41.41.4 using a threshold of $0.9$. A threshold of $0.01$ is sufficient for filtering out about $60\%$ of the features, increasing the mean generation success by around $0.4$ points. We compare this against a random baseline (green): filtering randomly sampled subsets of features of the same size does not lead to any significant improvements (results are average of 10 random samples per subset size).
The results in Section 3 suggest that features with high output scores occur predominantly in later layers. Figure 5 shows the generation success when filtering based on the output score, evaluated only on features from later layers of the model: 16–25 for Gemma2-2B and 24–41 for Gemma2-9B. Taking only features from later layers, Gemma2-2B and Gemma2-9B achieve generation success scores of about $0.8$. By considering only top-scoring features, the mean generation success increases to around 1.11.11.11.11.41.41.41.4 for both models. These results show that even within these later layers, the output score is a useful tool for filtering out features that lead to poor steering results.
Gemma2-9B-it
@@
Table 1: Results on the Concept500 dataset from AxBench on instruction-tuned Gemma-2-9B. (Top) Results when steering with SAEs after filtering out features with $S_{out}$ lower than different thresholds. (Bottom) Results reported by Wu et al. (2025). Bold indicates the best score, underline indicates the best score among representation-based methods. Grey indicates non-representation-based methods. After filtering based on output scores, SAEs achieve the best score among representation-based methods at L31, and reach 90.7% of the best method’s performance at L20.

### v2

@@ Abstract
Dana Arad1thanks: Work partially done during an internship at Amazon.   Aaron Mueller2   Yonatan Belinkov1 1Technion – Israel Institute of Technology  2Boston University danaarad@campus.technion.ac.ilamueller@bu.edu belinkov@technion.ac.il
@@ 1 Introduction
Sparse autoencoders (SAEs) have shown promise in extracting human-interpretable features from the hidden states of language models (LMs) bricken2023monosemanticity; scaling2024templeton. One appealing usage of SAEs is to enable fine-grained interventions such as generation steering o2024steering; durmus2024steering; marks2025sparse. However, selecting the right features for intervention is an open problem. Current approaches typically select features to steer based on their activation patterns, i.e., the input texts that most strongly activate a given feature interpretable2024huben. While input-based activations can reveal meaningful patterns, recent work highlights a critical limitation: a feature’s activations are not necessarily the same as its causal effect on the model’s output durmus2024steering; paulo2024automatically; gurarieh2025enhancing. As a result, the way features are selected can lead to suboptimal steering, reducing its consistency and reliability durmus2024steering.
@@
In this work, we formalize two distinct roles that features can play: input features, which capture patterns within the model’s input, and output features, whose main role is to directly influence the tokens the model generates. To find them, we propose input scores and output scores. First, we obtain a representative set of tokens for each feature by applying the logit lens to SAE weights; this projects the weights directly into the vocabulary space nostalgebraist; saelogitlens. We define input features as having high input scores—i.e., high overlap between their top-activating tokens and top logit lens tokens. We define output features as having high output scores—i.e., intervening on the feature increases the probability of its top logit lens tokens in the final output distribution. Notably, the input score can be computed in parallel for all features over a general dataset; the output score requires only one forward pass and no concept-specific data. We quantitatively show that these roles rarely co-occur and tend to emerge at different layers in the model. Specifically, features in earlier layers primarily act as detectors of input patterns, while features in later layers are more likely to drive the model’s outputs, consistent with prior analyses of LLM neuron functionality lad2024mstages; marks2025sparse.
@@
By calculating the input and output scores of features extracted from Gemma-2 (2B and 9B) team2024gemma, we show that features with high output scores are more effective for coherent and high-quality steering. This yields a practical feature selection method, illustrated in Figure˜1: starting with a (typically small) set of candidate features for steering, our scores are computed over this set to select a more effective subset. Figure˜2 demonstrates the difference when steering with input features vs. output features: steering with a feature that has a high input score but low output score fails to meaningfully influence the generation. In contrast, steering using a feature with high output score and low input score yields better steering, as well as more fluent and semantically coherent completions.
@@
We demonstrate the effectiveness of these insights on the recent AxBench wu2025axbench, a benchmark for evaluating steering methods. While AxBench found SAEs to be poor for steering, our feature selection results in a 2–3x improvement, causing SAE steering (an unsupervised method) to score significantly closer to supervised methods like LoRA xu2024lora.
@@
  • We propose a practical method for finding features effective for steering.
@@ 2.1 Sparse Autoencoders
Sparse Autoencoders (SAEs) were recently proposed as a method to address the problem of polysemanticity, where individual neurons entangle multiple unrelated concepts, and where a single concept may be distributed across many neurons bricken2023monosemanticity. Given a hidden representation $x\in\mathbb{R}^{n}$, an SAE consists of an encoder and a decoder, defined as:
@@
The encoder maps the latent $x$ into a higher-dimensional sparse vector $a(x)$, or $a$ for short, which we refer to as the activations. The decoder reconstructs $x$ from $a$ as a sparse linear combination of the learned features, given by the columns of $W_{\text{dec}}$. Sparsity and non-negativity of the activations are enforced through the non-linearity $\sigma$, often JumpReLU rajamanoharan2024jumprelu, and regularization.
@@ 2.2 The Logit Lens
The logit lens is a widely used interpretability tool for analyzing the hidden representations of language models nostalgebraist. Given a hidden state $x\in\mathbb{R}^{n}$ at any layer of the model, the logit lens passes $x$ through the final layer norm, LN, then projects $x$ onto the vocabulary space by applying the unembedding matrix $W_{\text{unembed}}\in\mathbb{R}^{n\times|\mathcal{V}|}$, where $\mathcal{V}$ is the model’s vocabulary. This produces a vector of predicted logits:
@@
Recent work has demonstrated that the logit lens can be applied not only to hidden representations, but also model weights dar2023analyzing, gradients katz2024backward, and even in multi-modal settings toker2024lens.
saelogitlens suggested applying the logit lens to SAE feature weights as a way to interpret their roles. To interpret an SAE feature $f_{i}$, corresponding to $W^{i}_{\text{dec}}$, the $i$-th column in the decoder matrix, we compute:
@@ 2.3 Steering LMs
We define steering as influencing the output of an LM towards a desired concept. Successful steering should maintain the quality and coherency of the generated text. In other words, we seek a minimal change to a model’s computation that adds or subtracts a concept’s influence. Steering can be done by various methods, recently using SAEs durmus2024steering; wu2025axbench.
Formally, given a model $M$ and a prompt $x$, we obtain a steered text $\tilde{y}$ by applying an intervention $\Phi(\cdot)$ on some intermediate representation $h$:
$\tilde{y}=M_{h\leftarrow\Phi(h)}(x)$ (5)
Similarly to scaling2024templeton, in order to steer an LM towards a concept $c$ encoded in an SAE feature $f_{i}$ at layer $l$, we define $\Phi$ as follows: first, we pass a prompt prefix $p$ through the model. At layer $l$, we pass the latent representation $x^{l}$ through the SAE encoder to obtain the activations vector, $a$. We record the max-activating feature, denoted $a_{\text{max}}$. Then, we obtain a new activation vector using steering factor $s$:
@@
Similarly to makeloveevaluating, to evaluate steering success we measure the generation success w.r.t. $\ell_{k}$ by calculating the number of appearances of any token in $\ell_{k}$ in the generated text. Given a set of sentences $S$:
$\text{Gen Success@k}(S)=\frac{\sum_{s\in S}{|\{t\in s\mid t\in\ell_{k}\}|}}{|S|}$ (7)
Additionally, we use perplexity measured using Gemma-2-9B to quantify the generation coherence of the entire generated text.
@@ 3.1 Input Features
In practice, we use the pre-computed activations from Neuronpedia neuronpedia to obtain the sentences $S$. We verify that $S$ has at least 20 sentences and take a maximum of 100 sentences per feature.
@@ 3.2 Output Features
Output features were first mentioned by paulo2024automatically as features whose effect on the model’s output can be easily explained in natural language. Natural language explanations predispose us to errors in both precision and recall (huang-etal-2023-rigorously); therefore, given a target concept, we quantify steering quality as consistency between the set of logit lens tokens and the set of tokens that the steering operation promotes.
@@
To measure the effect of a feature on the model’s output distribution, we perform an intervention during a forward pass and evaluate the change in the rank and probability assigned to tokens in $\ell$. We first use a neutral prompt ($x=$ ”In my experience,”) to obtain the model’s prior distribution over token ranks and probabilities. Then, we intervene on the feature’s activation value using a large steering factor (we use $10$), as in Equation˜6.
@@
where $V$ is the model’s vocabulary. If $x$ is a neutral prompt, then $S_{out}\propto P(\mathcal{M}_{h\leftarrow\Phi(h)})$, so we can compute $S_{out}$ quickly using a single forward pass by only computing the rank-weighted probability after the intervention. This score is robust to the specific choice of neutral prompt; see details in Appendix˜D.
@@ 3.3 Experimental Setup
We focus our analysis on Gemma-2 (2B and 9B) using the Gemma-Scope 16K SAEs team2024gemma; lieberum2024gemma, Llama-3.1 8B with Llama-Scope SAEs grattafiori2024llama; he2024llama, and Pythia-70m biderman2023pythia; interpretable2024huben. Our analysis spans 100 features randomly sampled from each layer. For Pythia-70m we limited our sampling to features with at least 10 recorded activations, since many features do not have any recorded activations on Neuronpedia.
(a) Gemma-2-2B.
(b) Gemma-2-9B.
Figure 3: Input and output scores across layers in Gemma-2-2B and Gemma-2-9B. The solid lines represent the median input score (blue) and output score (magenta), while the shaded regions denote the interquartile range (25th to 75th percentile), capturing the variability across features within each layer. Early layers are characterized by features with high input scores, while high output scores emerge in later layers.
@@ 3.4 Results
Figure˜3 shows the distribution of input and output scores across layers for Gemma-2-2B and Gemma-2-9B. In early layers (05050% of model depth), features tend to have high input scores and near-zero output scores, suggesting they are predominantly input-aligned. Later layers (6666100100% of model depth) show an opposite trend: input scores drop to near-zero, while output scores increase significantly. These later-layer features no longer reflect the tokens they are activated on, but instead align with the tokens they promote in the model’s output, indicating a shift toward output-aligned behavior. Interestingly, middle layers exhibit low scores for both metrics, suggesting that these features may play intermediate roles that are neither purely input-aligned nor strongly output-promoting.
For Llama-3.1 we do not observe any trend in early layers of the model (Appendix˜B), likely due to limitations of applying the logit lens to early layer representations nostalgebraist. At around 5050% of the model’s depth we begin to observe non-zero values for both scores, with low values of input scores and increasingly growing values of output scores as layers progress, as in the Gemma-2 results.
For Pythia, we observe a slightly different trend (Appendix˜B). While output score gradually increases around 5050% of the model’s depth as expected, the input score is mainly zero for most of the tested features. This may be due to the fact that this model is significantly smaller compared to the other models we examined ($70$ million parameters compared to 2299 billion), which may lead it to parse and encode information differently within its latent space.
Interestingly, and unlike Llama-3.1, early layers in both Gemma models and Pythia seem to be interpretable with the logit lens. We find that many features promote a coherent and human-understandable set of tokens, as reflected by high input scores as early as layers 0 and 1. (See Appendix˜C for examples.)
@@ 4.1 Experimental Setup
We use 50 prompt prefixes and generate up to 20 tokens, obtaining 50 generated texts for each feature (more details in Appendix˜E). For each feature we calculate the mean generation success across the generated texts, and filter out steering factors leading to generation success greater than $3$. Intuitively, the generation success measures the rate in which the model generates concept-related tokens. Based on early experiments, we find that 3 is a good upper value that balances steering and coherence. We choose the optimal steering factor as the one that maximizes $\frac{\text{Gen Success@20}}{\text{Perplexity}}$, where we normalize both metrics to a 011 range by dividing with the maximum value across all data samples.
@@ 4.2 Qualitative Results
Figure˜2 demonstrates steering with two features from Gemma-2-2B: one having a high input score and low output score (an input feature), and the other, an output feature, having a high output score and low input score. Steering using the input feature fails to meaningfully influence the generation. When using a high steering factor, this results in repeatedly generating tokens related to the feature’s activation, as if continuing from the word ”primary”. In contrast, steering using a feature with high output score and low input score using an optimal steering factor yields fluent and semantically coherent completions. Additional examples are shown in Appendix˜E.
(a) Gemma-2-2B.
(b) Gemma-2-9B.
Figure 5: Even in later layers of the model (16–25 for Gemma-2-2B and 24–41 for Gemma-2-9B), filtering features with low output scores increases mean generation success. Magenta: filtering by output scores. Green: filtering random sets of features of the same size.
@@ 4.3 Quantitative Results
Figure˜4 shows the mean generation success @ 20 of steered generations when filtering out features with output scores below varying thresholds (magenta) on Gemma-2-2B and Gemma-2-9B. As the threshold increases, performance improves steadily, indicating that features with higher output scores consistently lead to more successful steering. We observe an increase in the mean generation success score from around 0.50.60.5-0.6 for both models without any filtering, to 1.11.11.41.4 using a threshold of $0.9$. A threshold of $0.01$ is sufficient for filtering out about $60\%$ of the features, increasing the mean generation success by around $0.4$ points. We compare this against a random baseline (green): filtering randomly sampled subsets of features of the same size does not lead to any significant improvements (results are average of 10 random samples per subset size). Llama-3.1 and Pythia show similar trends; see Appendix˜B.
The results in Section˜3 suggest that features with high output scores occur predominantly in later layers. Figure˜5 shows the generation success when filtering based on the output score, evaluated only on features from later layers of the model: 16–25 for Gemma-2-2B and 24–41 for Gemma-2-9B. Taking only features from later layers, Gemma-2-2B and Gemma-2-9B achieve generation success scores of about $0.8$. By considering only top-scoring features, the mean generation success increases to around 1.11.11.41.4 for both models. These results show that even within these later layers, the output score is a useful tool for filtering out features that lead to poor steering results.
Gemma-2-9B-it
@@
Table 1: Results on the Concept500 dataset from AxBench on instruction-tuned Gemma-2-9B. (Top) Results when steering with SAEs after filtering out features with $S_{out}$ lower than different thresholds. (Bottom) Results reported by wu2025axbench. Bold indicates the best score, underline indicates the best score among representation-based methods. Grey indicates non-representation-based methods. After filtering based on output scores, SAEs achieve the best score among representation-based methods at L31, and reach 90.7% of the best method’s performance at L20.